OpenStudy (anonymous):

x+7/6 = 5/3 solve the proportion

OpenStudy (precal):

\[\frac{x+7}{6}=\frac{5}{3}\] cross multiply 3(x+7)=30 distribute 3 3x+21=30 -21 -21 3x=9 divide both sides by 3 x=3